# Spatial and Spectral Characterization of the Gravitational-wave Background with the PTA Optimal Statistic

### v1

@@ I Introduction
An alternative approach, and one which is most often used for determining detection significance, is the PTA optimal statistic (OS) [27]. This is a frequentist statistic originally developed to look for GWBs in ground-based gravitational-wave detectors [28]. In a PTA context, the OS was conceived to compute the amplitude and signal-to-noise ratio (S/N) of a GWB with a power-law strain spectrum, usually fixed to the expected shape given by a SMBHB population [27, 29]. Unlike the Bayesian approach, the OS focuses exclusively on the cross-correlated signal between pulsars, leveraging the distinctiveness of the HD curve and treating the auto-correlated GWB signal as a noise process that adds to intrinsic noise in the pulsars. It is also a much faster statistic to evaluate, which means that it is often the first approach carried out to provide detection statistics, GWB amplitude estimates, reconstructed correlation signatures, and cross-validation tests. Its speed also makes calibration of the OS S/N with large numbers of data augmentation operations (e.g., phase shifting [30] and sky scrambling [31]) more tractable.
@@
parameter estimates from a Bayesian CURN search by
iteratively running the OS with many parameter estimates.
More accurate in cases with significant intrinsic red noise.
@@ II.3 Parameter estimation
The problem, however, is that some PTA datasets are no longer in the weak-signal regime, as the GWB signal is becoming the dominant contribution to the total red noise in the lowest frequencies of many pulsars [5]. Revisiting our covariance matrix calculation,
@@ III A New Per-Frequency Optimal Statistic
The primary goal of this work is to further generalize the traditional OS to work for any GWB spectral shape, not just power-law behavior. We call the method presented here the Per-Frequency Optimal Statistic (PFOS). The traditional OS was constructed to measure the correlated amplitude and SNR of a power-law spectrum GWB signal, using a unit-amplitude power-law spectral template for the GWB, $\hat{\phi}=\phi/A^{2}_{\mathrm{gw}}$, where $\phi$ is represents the power-law PSD of the GWB in Equation 6. This broadband estimation condenses the correlated information from each frequency into a single, more sensitive measurement [27]. Recalling the rank-reduced forms in Equation 16 and Equation 17 for the cross-correlation estimator, we can change the spectrum which that estimator assumes by simply swapping $\hat{\phi}$ with another spectral template.
Here we change the spectral template to be entirely agnostic to the spectral shape. Since the traditional OS already deals in discrete Fourier frequencies in both the $\bm{X}_{a}$ and $\bm{Z}_{a}$ terms, we replace the unit-amplitude Fourier domain covariance matrix, $\hat{\phi}$, with a frequency-selector, $\tilde{\phi}(f_{k})$, which multiplies and adds the adjacent sine and cosine components of the two pulsars at the specified frequency. This frequency-selector will still have the same shape as $\phi$, ($2k_{\mathrm{max}}\times 2k_{\mathrm{max}}$) elements and still be diagonal, however, all of the elements of this diagonal matrix will be zero except for those which correspond to frequency $f_{k}$. This is best shown through examples,
@@
where $f_{k}$ represents the frequency to be analyzed (which must be a multiple of the lowest PTA frequency). However, swapping the spectral template from $\hat{\phi}$ to $\tilde{\phi}(f_{k})$ does introduce a new assumption that measurements at each frequency are independent of one another; we test this assumption in subsection V.3.
@@
Including $\Phi(f_{k})$ may seem counterintuitive, as it requires information on the spectrum we are currently trying to estimate. However, this is similar to the problem of incorporating $A^{2}_{\mathrm{gw}}$ as an assumed amplitude when using pulsar-pair covariance with the traditional OS. We employ the same strategy that was discussed in subsection II.3 where we use the CURN estimates of the GWB as the assumed spectral shape. As we will soon show, this scheme produces estimates that are consistent with injection values in our simulations while also allowing us to fit frequencies one at a time, and keeping our covariance matrices smaller and more stable. Details about effects resulting from the choice of a CURN model can be found in subsection V.2 and Section VI.
@@
We can make an additional simplification to the PFOS method if we assume that all the GW power comes from only the single frequency being analyzed, i.e., there is no measurable GWB signal in frequencies other than $f_{k}$. In this scheme, our “true” unit-amplitude spectral shape changes to be a Kronecker delta function at the desired frequency, such that $\Phi(f_{k})=\tilde{\phi}(f_{k})$ in Equation 28 and Equation 29.  [^3] This somewhat analogous to a GWB “radiometer” anisotropy search where we use a sky-position delta function for gravitational wave power. The narrowband-normalized PFOS instead uses a frequency specific delta function [43, 44]. We will refer to this form as the narrowband-normalized PFOS.
The narrowband-normalized PFOS has similar cross-correlation estimators and associated variances given by
@@
While the narrowband-normalized PFOS removes the need for an initial estimate of the spectral shape and simplifies the equations, it leads to slightly biased results in cases where there is a detectable signal in other frequencies. Based on our testing, this simplification does work well in cases where individual pulsars have similar observing cadence and length. However, as we will show in subsection V.2, using more realistic data shows worse performance than the broadband-normalized PFOS. Because of this, we recommend using the broadband-normalized PFOS unless the data set is within the weak-signal regime.
Finally, it is also possible to use the different flavors of PFOS to calculate SNRs at each frequency. Similar to the traditional OS, the PFOS’s null hypothesis is the absence inter-pulsar correlations. However, the null hypothesis should be interpreted differently depending on whether the broadband-normalized or narrowband-normalized PFOS is used. The broadband-normalized PFOS implies a null hypothesis in which one assumes that no measurable correlations exist in a particular frequency bin, but permits them to be present in other bins. This is less biased in cases where there are significant correlations in other frequency bins, making it possible to isolate the significance of a single frequency bin in a detectable spectrum. By contrast, the narrowband-normalized PFOS implies a null hypothesis in which one instead assumes that no frequency bins whatsoever have measurable correlations. Hence the narrowband-normalized PFOS null hypothesis is most useful in cases where there are not significant correlations present at any frequency, and as such we are trying to estimate the significance in a single frequency bin. These interpretations and the potential for other interpretations of the null hypothesis should be deliberated upon on a case-by-case basis, and should always be calibrated through empirical techniques [30, 42].
@@ V.2 Per-frequency PSD recovery
Shown in Figure 6 are P-P plots representing the first five frequencies of the $100$ simulations with the largest GWB amplitude of $A_{\mathrm{gw}}=10^{-14}$. These strong-signal regime simulations show the largest bias, and are representative of a worst-case scenario for the PFOS. We see the first four frequencies show mostly unbiased behaviors with an estimator mean that only slightly underestimates the true distribution while remaining well within the $3\sigma$ contours. The fifth frequency briefly departs this $3\sigma$ contour, indicating a slight underestimation of $S(f_{k})$ at that frequency. Higher frequencies (not shown) exhibit the same unbiased behavior as the lowest four frequencies. The slight bias in the fifth frequency is also present with weaker GWB amplitudes, but less pronounced. This isolation of a particular problematic frequency hints at a potential issue with the PTA configuration used in the simulations. This may result from the irregular timing cadences and gaps leaving poor sensitivity in this particular frequency, as more ideal simulations with even timing cadence show no bias at any frequency. More work must be done to determine exactly the cause of this, but since it does not appear to be an issue with the PFOS itself this remains beyond the scope of the current work.
We also test the narrowband-normalized PFOS which is more agnostic to the assumed spectral shape. The narrowband-normalized PFOS is slightly biased in situations in which there are detectable correlations at other frequencies of the simulated data. Figure 7 shows a P-P plot with the narrowband-normalized PFOS for the first five GW frequencies. Each frequency remains broadly consistent within the $3\sigma$ contours with only a minor deviation in the fourth frequency. We note however that the mean at every frequency is marginally overestimated. The uncertainty in the estimators for each frequency are also inflated due to the mis-modeled spectral shape. For these reasons, the more accurate and robust solution provided by the broadband-normalized PFOS are preferred whenever possible.
@@ V.3 Broadband spectrum recovery
In addition to agnostic spectral recovery, we also test the performance of fitting the PFOS’ PSD estimates with a parameterized spectral model. We opt for a power-law so that we can directly compare the PFOS model fitting to the traditional OS. As with the individual frequency tests, we source our CURN power-law estimates from Bayesian CURN analyses. Both the PFOS and traditional OS methods incorporate pulsar-pair covariance and, to provide a fair comparison, are provided with the same vectors of parameter estimates. Given the individual PSD estimates resulting from the PFOS, we carry out a least-squares fit for a power-law amplitude (as usual referenced to a frequency of $f={\rm yr}^{-1}$) with the spectral index $\gamma$ set to the value of the Bayesian CURN power-law spectral index parameter to exactly mimic the process of the traditional OS’ fitting. With noise marginalization and uncertainty sampling for both the PFOS and traditional OS, we find the percentiles at which the injected GWB amplitude lies in each of our $100$ simulations.
@@
Figure 8: Broadband P-P plots for a power-law amplitude of $A_{\mathrm{gw}}=10^{-14}$ for a linear power-law fit on the pair-covariant (broadband-normalized) PFOS and pair-covariant OS using uncertainty sampling and noise marginalization from a Bayesian CURN search. The shaded regions indicate the 1, 2, and 3 sigma contours from a perfectly unbiased estimator for these 100 simulations.
@@
The results for $A_{\mathrm{gw}}=10^{-14}$ are shown in Figure 8. As with the individual frequency analysis, the largest GWB amplitude produces the greatest deviations from an unbiased estimator, such that Figure 8 again represents a worst-case scenario. We see that the traditional OS with pulsar-pair covariance performs well and remains within the $3\sigma$ confidence contours at all percentiles, even in this strong signal regime. By contrast, the power-law fit on the PFOS analysis products produce minor deviations. There is a slight sag, especially with higher percentiles, which indicate that the PFOS’ spectral fit and uncertainty are both slightly too large. The fifth frequency problem shown in Figure 6 may be causing some of these discrepancies, as more idealized simulations with even cadences show even better agreement between the two methods.
@@
This test demonstrates that the PFOS’ ability to accurately measure the GWB PSD is sufficient to then use in a full spectral model fit. Despite the fact that we remove the power-law assumption of the traditional OS, this method can produce nearly equivalent results when both methods correctly account for pulsar-pair covariance.
@@ V.4 Stochastic background plus single GW signal
Despite the differences, nearly all PSD recoveries are performing well. Every frequency in the GWB-only case includes the injected value within the 95% credible region of the uncertainty sampled distributions. This remains the case for the GWB$+$CW injection for every frequency except the CW frequency. This frequency has the injected GWB+CW signal at just below the $3\sigma$ level; this isn’t too surprising, as neglecting the CW’s influence in the GWB leads to slightly poorer estimates of the total noise and the fact that the CW’s correlations will not match those of the HD.
@@ VI Discussion & Conclusions
Additionally, we found that in order to properly normalize the PFOS, it was important to introduce an estimate of the GWB spectral shape. This manifested as the unit-amplitude spectral shape at frequency $f$, $\Phi(f_{k})$, that is analogous to the unit-power-law spectral shape, $\hat{\phi}$, in the traditional OS. We opted to treat this spectral shape estimate in a similar manner to the pulsar-pair covariant OS where this estimate is dictated by the CURN model parameters used to construct the (broadband-normalized) PFOS. subsection V.2 and subsection V.3 shows that using this strategy results in consistent measurements with injected values. We also introduced a separate form of the PFOS called the narrowband-normalized PFOS, which is entirely agnostic to the spectral shape. This benefit comes at the cost of introducing slight bias when GWB-induced correlations are detectable in other frequencies. subsection V.2 showed that despite this purposeful mis-modeling, the narrowband-normalized PFOS still delivers valid individual-frequency measurements. However, we recommend that unless entirely independent frequency measurements are necessary, the more accurate broadband-normalized PFOS should be used.
@@
VII Acknowledgements

### v2

@@ I Introduction
An alternative approach, and one which is most often used for determining detection significance, is the PTA optimal statistic (OS) [27]. This is a frequentist statistic originally developed to look for GWBs in ground-based gravitational-wave detectors [28]. In a PTA context, the OS was conceived to compute the amplitude and signal-to-noise ratio (S/N) of a GWB with a power-law strain spectrum, usually fixed to the expected shape given by a SMBHB population [27, 29]. While other frequentist GWB detection techniques also exist and have been used in some recent publications (i.e., Jenet et al. [30] was used in Xu et al. [9]), the OS was shown to be both a minimum variance estimator within noise dominated data and an optimal detection statistic [27]. Unlike the Bayesian approach, the OS focuses exclusively on the cross-correlated signal between pulsars, leveraging the distinctiveness of the HD curve and treating the auto-correlated GWB signal as a noise process that adds to intrinsic noise in the pulsars. It is also a much faster statistic to evaluate, which means that it is often the first approach carried out to provide detection statistics, GWB amplitude estimates, reconstructed correlation signatures, and cross-validation tests. Its speed also makes calibration of the OS S/N with large numbers of data augmentation operations (e.g., phase shifting [31] and sky scrambling [32]) more tractable.
@@
parameter estimates from a Bayesian common uncorrelated
red noise search by iteratively running the OS with many
parameter estimates. More accurate in cases with significant
intrinsic red noise.
@@ II.3 Parameter estimation
The problem, however, is that some PTA datasets are no longer in the weak-signal regime, as the GWB signal is becoming the dominant contribution to the total red noise in the lowest frequencies of many pulsars [5]. This was first proposed in the LIGO case in Allen and Romano [28], and was revisited for the pulsar timing case in Allen and Romano [34]. The first evidence of this issue in realistic simulations was shown as a part of the NANOGrav 15 year publications [22] and as such, was included within the analysis in Figure 1 of the NANOGrav 15 year evidence paper [5]. The way to solve this problem is to recalculate our covariance matrix without any simplifying assumptions,
@@ III A New Per-Frequency Optimal Statistic
The primary goal of this work is to further generalize the traditional OS to work for any GWB spectral shape, not just power-law behavior. We call the method presented here the Per-Frequency Optimal Statistic (PFOS). The traditional OS was constructed to measure the correlated amplitude and SNR of a power-law spectrum GWB signal, using a unit-amplitude power-law spectral template for the GWB, $\hat{\phi}=\phi/A^{2}_{\mathrm{gw}}$, where $\phi$ represents the power-law PSD of the GWB in Equation 6. This broadband estimation condenses the correlated information from each frequency into a single, more sensitive measurement [27]. Recalling the rank-reduced forms in Equation 16 and Equation 17 for the cross-correlation estimator, we can change the spectrum which that estimator assumes by simply swapping $\hat{\phi}$ with another spectral template.
Here we change the spectral template to be entirely agnostic to the spectral shape. Since the traditional OS already deals in discrete Fourier frequencies in both the $\bm{X}_{a}$ and $\bm{Z}_{a}$ terms, we replace the unit-amplitude Fourier domain covariance matrix, $\hat{\phi}$, with a frequency-selector matrix, $\tilde{\phi}(f_{k})$, which multiplies and adds the adjacent sine and cosine components of the two pulsars at the specified frequency. This frequency-selector matrix will retain the same shape as $\phi$, ($2k_{\mathrm{max}}\times 2k_{\mathrm{max}}$) elements and still be diagonal, however, all of the elements of this diagonal matrix will be zero except for those which correspond to frequency $f_{k}$. This is best shown through examples,
@@
where $f_{k}$ represents the frequency to be analyzed (which must be a multiple of the lowest PTA frequency). However, swapping the spectral template from $\hat{\phi}$ to $\tilde{\phi}(f_{k})$ does introduce a new assumption that PTA GW frequency measurements are independent of one another; we test this assumption in subsection V.3.
@@
Including $\Phi(f_{k})$ may seem counterintuitive, as it requires information on the spectrum we are currently trying to estimate. However, this is similar to the problem of incorporating $A^{2}_{\mathrm{gw}}$ as an assumed amplitude when using pulsar-pair covariance with the traditional OS. We employ the same strategy that was discussed in subsection II.3 where we use the CURN estimates of the GWB as the assumed spectral shape. As we will soon show, this scheme produces estimates that are consistent with injection values in our simulations while also allowing us to fit frequencies one at a time and keeping our covariance matrices smaller and more stable. Details about effects resulting from the choice of a CURN model can be found in subsection V.2 and Section VI.
@@
III.1 Frequency-independent simplification
We can make an additional simplification to the PFOS method if we assume that all of our pulsars measure each GW frequency independently of each other frequency. This assumption is valid if every pulsar is timed with equal cadences and contains no data gaps. Under this additional assumption, the $Z$ matrices defined in equation Equation 15 become frequency block-diagonal. With a frequency block-diagonal matrix, the normalization calculated in subsection A2 become independent of the spectral shape matrix $\Phi(f_{k})$. We can then replace this spectral shape matrix with a much simpler frequency-selector matrix $\tilde{\phi}(f_{k})$. As this also re-normalizes the PFOS as a frequency-independent narrowband estimator, we will refer to this form as the narrowband-normalized PFOS.
Substituting this frequency-selector matrix into Equation 28 and Equation 29 leads us to write our cross-correlation estimators and associated variances as
@@ III.1 Frequency-independent simplification
This narrowband-normalized PFOS looks remarkably like the traditional OS’s estimators and variances in Equation 16 and Equation 17. This simplicity leads us to initially develop this form before a more rigorous calculation of the normalization was completed.
While the narrowband-normalized PFOS benefits from the lack of an estimated spectral shape, this simplicity comes at a cost. When making this assumption in data sets where these GW frequencies are not independent, the narrowband normalization ignores contributions from other frequencies, leading to overestimation of the GWB spectrum. These problems become more pronounced in larger signal scenarios.
Based on our testing, this simplification can work well in more ideal data sets where individual pulsars have similar observing cadence and length. However, as we will show in subsection V.2, using more realistic data shows worse performance than the broadband-normalized PFOS. Additionally, even in data sets where the narrowband-normalized PFOS simplification can be applied, the broadband-normalized PFOS will result in identical answers. For these reasons, while we elect to show this as possible simplification, we strongly recommend using the broadband-normalized PFOS instead.
III.2 Signal-to-Noise ratio interpretations
Finally, it is also possible to use the different flavors of PFOS to calculate SNRs at each frequency. Similar to the traditional OS, the PFOS’s null hypothesis is the absence inter-pulsar correlations. However, the null hypothesis should be interpreted differently depending on whether the broadband-normalized or narrowband-normalized PFOS is used.
The broadband-normalized PFOS implies a null hypothesis in which one assumes that no measurable correlations exist in a particular frequency bin, but permits them to be present in other bins. This is less biased in cases where there are significant correlations in other frequency bins, making it possible to isolate the significance of a single frequency bin in a detectable spectrum.
By contrast, the narrowband-normalized PFOS implies a null hypothesis in which one instead assumes that no frequency bins whatsoever have measurable correlations. Hence the narrowband-normalized PFOS null hypothesis is most useful in cases where there are not significant correlations present at any frequency, and as such we are trying to estimate the significance in a single frequency bin. These interpretations and the potential for other interpretations of the null hypothesis should be deliberated upon on a case-by-case basis, and should always be calibrated through empirical techniques [31, 43].
@@ V.2 Per-frequency PSD recovery
Shown in Figure 6 are P-P plots representing the first five frequencies of the $100$ simulations with the largest GWB amplitude of $A_{\mathrm{gw}}=10^{-14}$. These strong-signal regime simulations show the largest bias compared with the weak and intermediate signal regimes, and are representative of a worst-case scenario for the PFOS. This figure shows a general trend of underestimation for all frequencies and percentiles. This is an indication of an underestimated mean [11]. We also see that while the first four frequencies remain within the $3\sigma$ contours, the fifth frequency briefly departs this contour at the $50^{\rm th}$ percentile. Higher frequencies (not shown) exhibit the same slightly underestimated but relatively unbiased behavior as the lowest four frequencies. The increase in bias at the fifth frequency around the $50^{\rm th}$ percentile is also present with weaker GWB amplitudes but are less pronounced. Notably, in more ideal simulations the PFOS recovery has no such problems with any frequencies. This isolation of a single problematic frequency in realistic data may indicate an issue with the PFOS recovery with some PTA configurations. Differences between these and more ideal simulations including irregular timing cadences, timing gaps, and non-uniform timing baselines, may cause a break down in the frequency independence of the PFOS. More work must be done to determine exactly when these discrepancies are problematic. However, as subsection V.3 will soon show, the effects of this bias are small enough to create consistent measurements with the traditional pipeline. It is expected that the PFOS will be at least somewhat biased, as to be completely unbiased would require knowledge of the exact shape of the GWB spectrum beforehand.
We also test the narrowband-normalized PFOS which is agnostic to the assumed spectral shape. Since the narrowband-normalized PFOS as constructed is mathematically biased in non-ideal data, we expect this method to perform worse overall. Figure 7 shows a P-P plot of the first 5 GW frequencies of the same simulations. This time analyzed using the narrowband-normalized PFOS. Surprisingly, each frequency remains broadly consistent within the $3\sigma$ contours with only a minor deviation in the fourth frequency. We note however that the mean at every frequency is systematically overestimated. This is expected behavior as the smaller overall weighting of the normalization shifts these estimates to higher values. For these reasons, the more accurate and robust solutions provided by the broadband-normalized PFOS are preferred.
@@ V.3 Broadband spectrum recovery
In addition to agnostic spectral recovery, we also test the performance of fitting the PFOS’ PSD estimates with a parameterized spectral model. We opt for a power-law so that we can directly compare the PFOS model fitting to the traditional OS. As with the individual frequency tests, we source our CURN power-law estimates from Bayesian CURN analyses. All three methods incorporate pulsar-pair covariance and, to provide a fair comparison, are provided with the same vectors of parameter estimates. Given all 10 frequencies of the individual PSD estimates resulting from the two different PFOS methods, we carry out a least-squares fit for a power-law amplitude (referenced to a frequency of $f={\rm yr}^{-1}$) with the spectral index $\gamma$ set to the value found in the Bayesian CURN power-law spectral index parameter. This exactly mimics the process of the traditional OS’s fitting. With noise marginalization and uncertainty sampling for both PFOS methods and traditional OS, we find the percentiles at which the injected GWB amplitude lies in each of our $100$ simulations.
@@
Figure 8: Broadband P-P plots for a power-law amplitude of $A_{\mathrm{gw}}=10^{-14}$ for a linear power-law fit on the pair-covariant broadband-normalized PFOS (b-PFOS), narrowband-normalized PFOS (n-PFOS), and pair-covariant OS using uncertainty sampling and noise marginalization from a Bayesian CURN search. The shaded regions indicate the 1, 2, and 3 sigma contours from a perfectly unbiased estimator for these 100 simulations.
@@
The results for $A_{\mathrm{gw}}=10^{-14}$ are shown in Figure 8. As with the individual frequency analysis, the largest GWB amplitude produces the greatest deviations from an unbiased estimator, such that Figure 8 again represents a worst-case scenario. We see that the traditional OS with pulsar-pair covariance, labeled as “OS”, performs well and remains within the $3\sigma$ confidence contours at all percentiles, even in this strong signal regime, though it does have some minor bias at lower percentiles. By contrast, the power-law fit on the broadband-normalized PFOS analysis, labeled as “PL fit b-PFOS”, contain other minor deviations. There is a slight sag, especially with higher percentiles but surprisingly does better than the traditional OS in lower percentiles. The fifth frequency problem and general underestimation shown in Figure 6 may be causing some of the discrepancies in higher percentiles, as more idealized simulations with even cadences show even better agreement between the two methods. Regardless, these P-P plots still show remarkable agreement despite the high strength of the GWB used.
Unlike the previous two cases, the narrowband-normalized PFOS, labeled as “PL fit n-PFOS”, vastly overestimates the injection at all percentiles while the uncertainty estimates were nearly identical to the broadband-normalized. This is a result of using the individual frequency overestimates which are then compounded into a worse broadband spectral fit. This further reinforces our recommendations to use the broadband-normalized PFOS instead.
@@
These tests demonstrate that the PFOS’s ability to accurately measure the GWB PSD is sufficient to then use in a full spectral model fit. Despite the fact that we remove both the power-law assumption of the traditional OS and any frequency covariance information, the broadband-normalized PFOS can produce nearly equivalent results when both methods correctly account for pulsar-pair covariance.
@@ V.4 Stochastic background plus single GW signal
Despite the differences, nearly all PSD recoveries are performing well. Every frequency in the GWB-only case includes the injected value within the 95% credible region of the uncertainty sampled distributions. This remains the case for the GWB$+$CW injection for every frequency except the CW frequency. This frequency has the injected GWB+CW signal at just below the $3\sigma$ level; this isn’t too surprising, as neglecting the contribution of the CW to the overall PSD leads to worse estimates of the total noise as well as the fact that the CW’s correlations will not match those of the HD.
@@ VI Discussion & Conclusions
Additionally, we found that in order to properly normalize the PFOS, it was important to introduce an estimate of the GWB spectral shape. This manifested as the unit-amplitude spectral shape at frequency $f$, $\Phi(f_{k})$, that is analogous to the unit-power-law spectral shape, $\hat{\phi}$, in the traditional OS. We opted to treat this spectral shape estimate in a similar manner to the pulsar-pair covariant OS where this estimate is dictated by the CURN model parameters used to construct the (broadband-normalized) PFOS. subsection V.2 and subsection V.3 shows that using this strategy results in consistent measurements with injected values. We also introduced a separate form of the PFOS called the narrowband-normalized PFOS, which is entirely agnostic to the spectral shape. This benefit comes at the cost of introducing bias when individual pulsars contain frequency covariance in their measurements of their noise. subsection V.2 showed that despite this purposeful mis-modeling, the narrowband-normalized PFOS still delivers only slightly overestimated individual-frequency measurements. However, due to the systematic errors that are encountered when fitting a broad-spectrum model, it is preferred to use the far greater accuracy attained by the broadband-normalized PFOS.
@@
VII Software
The software used to implement these and other methods have been included in a comprehensive analysis python software package called DEFIANT. This software was made to incorporate the various expansions of the PTA optimal statistic, including the methods described in this paper, through modular, combinatory components. These components also include methods not featured including the multi-component optimal statistic [36] and vastly improve on the performance of previous implementations with the inclusion of the pair covariance generalization [35, 22].
This software makes frequent use of the ENTERPRISE[50] pulsar timing array software package for many of the needed matrix products and quantities.
VIII Acknowledgements